### v1

# The systemic recoil velocity distribution and the scale height of isolated millisecond pulsar systems: Implications on neutron star retention fractions in star clusters

Abstract
Aims. The systemic recoil velocity ($v_{\mathrm{sys}}$) distribution of millisecond pulsars (MSPs) is essential for understanding the MSP formation channel(s) and for estimating the retention fractions of MSPs in star clusters, which can potentially be determined with precise astrometry of MSPs. However, the determination is complicated by MSPs’ long-term dynamic evolution and the scarcity of radial velocity measurements. This work looks to overcoming the complexity and deriving the MSP $v_{\mathrm{sys}}$ distribution from high-precision astrometric measurements.
Methods. We compiled 65 isolated (or field) MSP systems (including lone MSPs) that are well astrometrically determined, and calculated their transverse peculiar (or space) velocities $\bm{v}_{\perp}$ and Galactic heights $z$. Assuming that the Galactic-longitude components $v_{\mathrm{l}}$ of $\bm{v}_{\perp}$ are statistically stable over time (the “stable-$v_{\mathrm{l}}$” assumption), we approached the distribution of the $v_{\mathrm{l}}$ components of $\bm{v}_{\mathrm{sys}}$ by the observed $v_{\mathrm{l}}$ sample. Under the “isotropic-$\bm{v}_{\mathrm{sys}}$” assumption that $\bm{v}_{\mathrm{sys}}$ directions are uniformly distributed, we derived the MSP $v_{\mathrm{sys}}$ distribution from the distribution of the $v_{\mathrm{l}}$ component of $\mathrm{v}_{\mathrm{sys}}$. Based on the derived $v_{\mathrm{sys}}$ distribution, we tested the “stable-$v_{\mathrm{l}}$” assumption with dynamical population synthesis (DPS). Additionally, by matching the observed $z$ and the Galactic-latitude components $v_{\mathrm{b}}$ of $\bm{v}_{\perp}$ to the DPS counterparts, we estimated the initial and the current Galaxy-wide scale heights of isolated MSP systems.
Results. We find that the observed $v_{\mathrm{l}}$ sample can be well described by a linear combination of three normal distributions. Accordingly, the MSP $v_{\mathrm{sys}}$ distribution can be approximated by a linear combination of three Maxwellian components. Our DPS analysis verified the “stable-$v_{\mathrm{l}}$” assumption in the parameter space of this work, and estimated the initial and the current Galaxy-wide scale heights of isolated MSP systems to be about 0.32 kpc and 0.66 kpc, respectively.
[…]
Key Words.:
stars: kinematics and dynamics – pulsars: general – stars: neutron – astrometry – parallaxes – proper motions
1 Introduction
Compact stars are normally the dense remnants of stars that have exhausted their nuclear fuel, resulting in objects such as white dwarfs (WDs), neutron stars (NSs), and black holes (BHs). Distinct formation mechanism(s) have been proposed for different types of compact stars. For example, stellar BHs are believed to form mainly through either direct collapse (of massive stars) or a delayed collapse (e.g. Mirabel, 2017). In the latter scenario, a progenitor star’s core, following a supernova explosion, collapses into a proto-NS first, then succumbs to further collapse into a BH due to fallback accretion (e.g. Zhang et al., 2008). As another major form of compact stars, NSs can be born with core-collapse supernovae (CCSNe), while having alternative formation channels such as accretion-induced collapse (AIC; e.g. Tauris et al., 2013) and WD-WD mergers (e.g. Levan et al., 2006). Moreover, NSs encompass various subgroups such as magnetars and millisecond pulsar (MSP) systems, with each NS subgroup having its primary and alternative formation theories (e.g. Ding, 2022).
As distinct formation channels of compact stars may lead to different natal kicks imparted to the compact star, kinematic investigation of compact stars holds a key to discriminating between different formation channels. This investigation can be realized with precise astrometry using space-based telescopes operating at optical/infrared or through very long baseline interferometry (VLBI) observations made at radio, which has been pursued in the study of stellar BHs (e.g. Mirabel et al., 2002; Willems et al., 2005; Atri et al., 2019) and various NS subgroups (e.g. Tendulkar et al., 2013; Gonzalez et al., 2011; Ding et al., 2023, 2024b, 2024c; Disberg et al., 2024b).
In practice, the birth velocity and position of a compact star or a compact star system (i.e., a gravitationally bound system of $\lesssim 3$ stars hosting a compact star) can be derived with the knowledge of 1) the observed 3-dimensional (3D) position, 2) the observed 3D velocity and 3) the age of the system, by integrating the motion backward over time assuming a model of Galactic potential. However, the age of a compact star system is usually not well constrained, therefore requiring additional assumption about the birth site (e.g. thin Galactic disk) of the compact star system (e.g. Fragos et al., 2009; Atri et al., 2019). Furthermore, most compact star systems, especially pulsar systems, do not have the information of radial velocities $v_{\mathrm{r}}$. As a result, the investigation of the systemic recoil velocities $\bm{v}_{\mathrm{sys}}$ (the recoil velocities of lone compact stars or stellar systems hosting compact stars) of pulsar systems has been mainly limited to kinematically young ($\lesssim 10$ Myr old) systems (e.g. Cordes & Chernoff, 1998; Kramer et al., 2003; Lorimer et al., 2006; Verbunt et al., 2017; Igoshev, 2020; Ding et al., 2024c) in order to minimize the dynamic evolution (of the compact star systems) through the Galactic potential.
On the other hand, for any specific subgroup of kinematically old ($\gtrsim 40$ Myr) compact star systems that do not have $v_{\mathrm{r}}$ information, the observed transverse peculiar velocities $\textit{{v}}_{\perp}$ (also referred to as transverse space velocity, i.e., velocity component tangential to the line of sight, with respect to the standard of rest of the neighborhood) is expected to deviate from the $v_{\mathrm{sys}}$ component tangential to the sightline (e.g. Disberg et al., 2024a). Despite the deviation, it has been proposed by Faucher-Giguere & Kaspi (2006) that the Galactic-longitude component $v_{\mathrm{l}}$ of $\bm{v}_{\perp}$ is statistically stable over time, which serves as a critical assumption that can link $v_{\mathrm{l}}$ to the $v_{\mathrm{sys}}$ distribution. For this assumption to be validated in the parameter space of interest, dynamical population synthesis (DPS) is required, which simulates the motion of stellar objects through the Galactic potential assuming initial states (i.e., positions and velocities) (e.g. Cordes & Chernoff, 1997).
[…]
Up till now, a well defined MSP $v_{\mathrm{sys}}$ distribution remains largely absent, mainly due to the aforementioned technical challenges and relatively large astrometric uncertainties. This absence hinders the studies of MSP formation channels, and results in weak constraints on the NS retention fractions in star clusters. Recent years have seen a major increase in the number of well astrometrically determined MSPs (e.g. Ding et al., 2023; Shamohammadi et al., 2024), which promises a refined characterization of the MSP $v_{\mathrm{sys}}$ distribution. In this paper, we determine the MSP $v_{\mathrm{sys}}$ distribution with the observed $v_{\mathrm{l}}$ calculated for 65 well astrometrically determined MSPs, then discuss its implications for the NS retention fractions in star clusters. Throughout this paper, uncertainties are stated at 68% confidence.
2 The observed transverse peculiar velocities and Galactic heights of millisecond pulsars
The establishment of the observed $\bm{v}_{\perp}$ distribution requires precise determination of both proper motions and distances for the MSPs, especially as MSP $v_{\perp}$ are much lower than those of normal pulsars (Hobbs et al., 2005; Gonzalez et al., 2011). Applying low-precision astrometric results to MSP kinematics studies would not only shift the average $v_{\perp}$ to the higher end, but also reduce the ability to resolve the potential multiple modes in the $\bm{v}_{\perp}$ distribution. Compared to other methods, the determination of trigonometric parallaxes through VLBI astrometry (e.g. Ding et al., 2023), pulsar timing (e.g. Perera et al., 2019; Shamohammadi et al., 2024), or Gaia astrometry (e.g. Jennings et al., 2018; Moran et al., 2023), offers the most precise and reliable distances for MSPs. Recent releases of the astrometric results by Ding et al. (2023) and Shamohammadi et al. (2024) have significantly enriched the sample of well astrometrically determined MSPs, which offers a rare opportunity to make statistically significant investigation of the MSP kinematics based purely on well astrometrically determined MSPs.
We searched for viable MSPs within the PSRCAT Catalogue [^1] https://www.atnf.csiro.au/research/pulsar/psrcat/ (Manchester et al., 2005). Due to relatively frequent dynamic interactions inside globular clusters, the $\bm{v}_{\perp}$ of globular cluster MSPs are likely not reflective of their $v_{\mathrm{sys}}$. With globular cluster MSPs excluded from the study, we reached 65 isolated MSP systems (sometimes referred to as field MSP systems) having significant ($>3\,\sigma$) proper motion and parallax determinations (see Table 2). The details of the sample selection are described in Appendix A.
Following the methods detailed in Section 6 of Ding et al. (2023), we derived the distances $d$ and the $\bm{v}_{\perp}$ of the 65 MSPs from their parallaxes and proper motions. From the $d$, we calculated the Galactic heights $z\equiv d\sin{b}$ for the 65 MSPs, where $b$ refer to Galactic latitudes. The derived $d$, $\bm{v}_{\perp}$ and $z$ are summarized in Table 2, where the columns related to $\bm{v}_{\perp}$ include the $\bm{v}_{\perp}$ magnitudes $v_{\perp}$, the Galactic-longitude component $v_{\mathrm{l}}$ and the Galactic-latitude component $v_{\mathrm{b}}$ of $\bm{v}_{\perp}$. On top of $v_{\mathrm{l}}$ and $v_{\mathrm{b}}$, $z$ serves as an additional indicator of MSP kinematics, as the average level of $|z|$ expected to grow with higher $v_{\mathrm{sys}}$. Therefore, this work includes the studies of three samples of measurements: the $v_{\mathrm{l}}$, the $v_{\mathrm{b}}$, and the $z$ samples.
Based on the calculations of the Pearson coefficients, no linear correlation is found between any two of the three samples. Accordingly, the three samples are treated independently. We conducted Wilcoxon signed-rank test on each of the three samples of 65 measurements, and confirmed no indication of asymmetry about 0 in any sample. Accordingly, in the following analysis, we assumed that the samples of the observed $v_{\mathrm{l}}$, $v_{\mathrm{b}}$ and $z$ are symmetric about 0.
To 1) visualize the three samples of measurements, and 2) investigate the underlying probability density function (PDF) for each sample, we assumed each measurement follows a split normal distribution, simulated 10,000 values per measurement, and concatenated the simulations into a chain of 650,000 values for each of $v_{\mathrm{l}}$, $v_{\mathrm{b}}$ and $z$. Despite the simulated sample of 650,000 following essentially a multi-modal distribution, the visualization is expected to characterize where the measurements overlap. The resultant histograms are displayed in Figure 2. According to Table 2 and Figure 2, most of the 65 MSPs are within the scale height of the Galactic thick disk (i.e., $\sim 1$ kpc).
Figure 2: The upper-left, lower-left and lower-right panels display the histograms of Galactic heights $z$, the Galactic-longitude component $v_{\mathrm{l}}$ and the Galactic-latitude component $v_{\mathrm{b}}$ of transverse peculiar velocities, respectively. Each histogram is concatenated from 10000 simulations drawn from the assumed split normal distributions for the 65 measurements (of $z$, $v_{\mathrm{l}}$ or $v_{\mathrm{b}}$; see Table 2). The gray dashed curves show the normalized count smoothed with kernel density estimation (using scipy.stats.Gaussian_kde). The best-fit Cauchy, Laplace and three-component normal distributions are plotted in blue and green dashed curves. Overlaid are the dynamical population synthesis results (labeled as “evolved”) that best match the observed $z$, $v_{\mathrm{l}}$ and $v_{\mathrm{b}}$, based on the determined initial MSP scale height $\zeta_{0}=0.32$ kpc. The corresponding initial distributions are also provided except for $v_{\mathrm{l}}$, as the initial $v_{\mathrm{l}}$ distribution adopted in the dynamical population synthesis is identical to the best-fit three-component normal distribution (a linear combination of three normal distributions).
Following the method detailed in Section 7.2.3 of Ding et al. (2024b), we ran Hartigan’s dip tests (Hartigan & Hartigan, 1985) on the three samples (of 65 measurements), and found no indication of multi-modality in any of the samples. Should the $v_{\mathrm{l}}$ sample truly follow the multi-modal distribution used in the visualization, about 200 and 260 additional $v_{\mathrm{l}}$ measurements would be needed to rule out uni-modality of the $|v_{\mathrm{l}}|$ sample at 90% and 95% confidence, respectively, according to our further simulations.
3 The systemic recoil velocity distribution
We followed Faucher-Giguere & Kaspi (2006) to approach the $v_{\mathrm{sys}}$ distribution with the observed $v_{\mathrm{l}}$ sample (see Table 2), assuming that $v_{\mathrm{l}}$ stays statistically stable over the long ($\lesssim 10$ Gyr) journeys of MSPs. The validation of this “stable-$v_{\mathrm{l}}$” assumption is detailed in Section 4. We determined the $v_{\mathrm{sys}}$ distribution in mainly three steps.
In the first step, we identified PDF candidates based on two simplifying assumptions. Given no indication that the observed $v_{\mathrm{l}}$ sample is multi-modal or asymmetric about zero (see Section 2), we assumed that the underlying PDF of the observed $v_{\mathrm{l}}$ sample is both uni-modal and symmetric about zero. Basic PDFs that meet these two requirements include (but are not limited to) normal distributions, Laplace distributions, and Cauchy distributions. Additionally, linear combinations of these basic distributions also serve as useful candidates. In particular, a linear combination of normal distributions with different scales (hereafter referred to as a multi-component normal distribution) would conveniently correspond to a multi-component Maxwell distribution for $v_{\mathrm{sys}}$ (assuming that $\bm{v}_{\mathrm{sys}}$ is equally possible in all directions for kinematically old pulsar systems).
In the second step, we fit the parameter(s) of each candidate PDF using scipy.optimize.curve_fit (Virtanen et al., 2020), for the PDF to approximate the curve (see Figure 2) smoothed from the normalized histogram of the measurement-based simulated $v_{\mathrm{l}}$ sample of 650,000 (see Section 2), where the smoothing was conducted with kernel density estimation using scipy.stats.Gaussian_kde (Virtanen et al., 2020). After discarding candidates with poor fitting results, we identified two groups of prime PDF candidates (that well match the normalized $v_{\mathrm{l}}$ histogram), with one group being linear combinations of Cauchy distributions and the other group being linear combinations of the normal distributions. The prime candidate PDFs are described in detail in Appendix B.1.
In the final step, we selected the PDF for the observed $v_{\mathrm{l}}$ sample by the method detailed in Appendix B.2. We conclude that the observed $v_{\mathrm{l}}$ can be well described by a Cauchy distribution with scale of 51.8 $\mathrm{km\leavevmode\nobreak\ s^{-1}}$ (i.e., Equation 8). Despite a three-component normal distribution being more consistent with the Cauchy distribution (see Figure 2), the more complex distribution has four more free parameters, therefore being less favorable. On the other hand, the three-component normal distribution, described by Equation 3 and Table 4, not only serves as a good approximation to the underlying PDF of the observed $v_{\mathrm{l}}$ sample, but also provides a rare analytical solution for the $\bm{v}_{\mathrm{sys}}$ magnitude $v_{\mathrm{sys}}$: when assuming the $\bm{v}_{\mathrm{sys}}$ distribution is isotropic for kinematically old pulsar systems, $v_{\mathrm{sys}}$ is expected to follow a three-component Maxwell distribution (i.e., a weighted sum of three Maxwell distributions) that share the same parameters with the best-fit three-component normal distribution (see Table 4). Namely, under the “isotropic-$\bm{v}_{\mathrm{sys}}$” assumption, $v_{\mathrm{sys}}$ should approximately follow the PDF
$\begin{split}p=&\sqrt{\frac{2}{\pi}}\Biggl{\{}0.498\cdot\frac{{v_{\mathrm{sys}}}^{2}}{45.5^{3}}\cdot\exp{\left[-\frac{1}{2}\left(\frac{v_{\mathrm{sys}}}{45.5}\right)^{2}\right]}\\ &\quad+0.447\cdot\frac{{v_{\mathrm{sys}}}^{2}}{110.3^{3}}\cdot\exp{\left[-\frac{1}{2}\left(\frac{v_{\mathrm{sys}}}{110.3}\right)^{2}\right]}\\ &\quad+0.055\cdot\frac{{v_{\mathrm{sys}}}^{2}}{448.8^{3}}\cdot\exp{\left[-\frac{1}{2}\left(\frac{v_{\mathrm{sys}}}{448.8}\right)^{2}\right]}\Biggl{\}}\,\,,\end{split}$ (1)
where the unit of $p$ and $v_{\mathrm{sys}}$ is $\mathrm{s\leavevmode\nobreak\ km^{-1}}$ and $\mathrm{km\leavevmode\nobreak\ s^{-1}}$, respectively. In contrast, no analytical solution of the $v_{\mathrm{sys}}$ PDF can be solved from the best-fit Cauchy distribution of $v_{\mathrm{l}}$. Therefore, we adopt Equation 1 as the $v_{\mathrm{sys}}$ PDF in the following analysis, which is illustrated along with its cumulative distribution function (CDF) in Figure 4.
Regarding the use of Equation 1, we like to clarify two points. Firstly, despite being a three-component Maxwell distribution, Equation 1 is still a uni-modal distribution by definition (see Figure 4). Secondly, though approximating the $v_{\mathrm{sys}}$ PDF by Equation 1, we do not suggest that isolated MSP systems have multiple formation channels. To further the investigation of formation channels of isolated MSP systems, population synthesis studies based on various MSP formation channels are needed to generate the theoretical $v_{\mathrm{sys}}$ PDF templates that can be compared with our observation-based counterpart.
4 Dynamical population synthesis
When deriving the $v_{\mathrm{sys}}$ PDF, we assumed $v_{\mathrm{l}}$ to be statistically stable over time (see Section 3). We tested this “stable-$v_{\mathrm{l}}$” assumption using DPS that simulates the dynamics of isolated MSP systems in the Galaxy. Specifically, we like to examine if the current (or evolved) $v_{\mathrm{l}}$ distribution stays roughly the same as the initial $v_{\mathrm{l}}$ distribution. We used Equation 1 as the MSP $v_{\mathrm{sys}}$ PDF in the DPS analysis. The MWPotential2014 model, calculated and compiled into the galpy package by Bovy (2015), was adopted as the Galactic potential used in the DPS. Apart from the $v_{\mathrm{sys}}$ PDF and the Galactic potential, two other ingredients are essential the DPS analysis — the MSP age distribution, and the initial spatial distribution of isolated MSP systems in the Galaxy. The former ingredient is described in Appendix C.1, while the latter is introduced as follows.
4.1 The initial Galactic spatial distribution
The initial Galactic distribution of isolated MSP systems consists of two parts — the birth radial distribution (i.e., the spatial distribution projected to the Galactic plane) and the birth Galactic height distribution (hereafter referred to as the $z_{\mathrm{i}}$ distribution). For the former part, we adopted the pulsar birth radial distribution by Xie et al. (2024), which assumes that Galactic isolated radio pulsars are born in the spiral arms.
On the other hand, the $z_{\mathrm{i}}$ distribution of isolated MSP systems is largely unknown. As we found that the observed $z$ sample (see Table 2) can be well described by a Laplace distribution with the scale height $\zeta^{\mathrm{local}}_{\mathrm{obs}}=0.52$ kpc (see Figure 2 and Appendix C.2.1), we assumed that $z_{\mathrm{i}}$ of Galaxy-wide isolated MSP systems also follow a Laplace distribution with the scale height of $\zeta_{0}$. In DPS, we realized that both the evolved $v_{\mathrm{b}}$ and the evolved $z$ change with $\zeta_{0}$ (see Appendix C.2.2). Therefore, by matching the evolved $v_{\mathrm{b}}$ and the evolved $z$ to the observed counterparts, one can determine $\zeta_{0}$. The DPS procedure that determines the evolved states (including the evolved positions and the evolved velocities) is described in Section 4.2, while the details of the $\zeta_{0}$ determination are provided in Appendix C.2.2. We find that the DPS with $\zeta_{0}=0.32$ kpc renders evolved $v_{\mathrm{b}}$ and evolved $z$ samples that well match the observed counterparts (see Figure 10). Accordingly, we adopt $\zeta_{0}=0.32$ kpc for the Galaxy wide $z_{\mathrm{i}}$ distribution of isolated MSP systems.
4.2 The procedure of dynamical population synthesis
In DPS, we first simulated 10,000 initial positions based on the initial spatial distribution described in Section 4.1. At each position, we assumed the stellar progenitor of the MSP followed circular motion (about the axis of rotation of the Galaxy) specified by the rotation curve of the MWPotential2014 Galactic potential model. $\bm{v}_{\mathrm{sys}}$ were simulated based on the $v_{\mathrm{sys}}$ PDF (i.e., Equation 1) and the “isotropic-$\bm{v}_{\mathrm{sys}}$” assumption (see Section 3). By adding the simulated $\bm{v}_{\mathrm{sys}}$ to the position-based circular velocities, we obtained the simulated MSP velocities with respect to the Galactic centre. Based on the simulated initial positions and velocities, we calculated the simulated initial $v_{\mathrm{l}}$, $v_{\mathrm{b}}$, and $z_{\mathrm{i}}$ in the same way as the calculation of the observed counterparts. The histograms of the initial $v_{\mathrm{b}}$ and $z_{\mathrm{i}}$ samples were smoothed with kernel density estimation (using scipy.stats.Gaussian_kde), and overlaid to Figure 2. As the three-component normal distribution we adopt as the initial $v_{\mathrm{l}}$ distribution is already displayed in Figure 2, we do not overlay the smoothed histogram of the initial $v_{\mathrm{l}}$ sample.
Subsequently, we performed orbit integrations with galpy to obtain the evolved positions and velocities after the span of each age in the sample of 13,005 MSP ages described in Appendix C.1. In this way, we compiled $10,000\times 13,005=130$ million evolved states (i.e., positions and velocities for different MSP ages) of simulated MSPs, from which $\sim 4$% evolved states outside the edge of the Galaxy (i.e., beyond 292 kpc from the Galactic centre, Deason et al., 2020) were discarded. To take into account the selection effect that astrometrically well determined pulsars are predominantly located within $\lesssim 8$ kpc, we removed evolved states that are beyond the 8 kpc distance. Additionally, given that the observed MSPs compiled in this work are $\lesssim 2.5$ kpc from the Galactic plane (see Figure 2), we discarded the evolved states with $|z_{\mathrm{e}}|>2.5$ kpc, where $z_{\mathrm{e}}$ refer to the evolved Galactic heights. The remaining evolved states are $\sim 37$ million in number, which amount to $\sim 30$% of the Galaxy-wide evolved states.
4.3 Results of dynamical population synthesis
From the remaining evolved states, we calculated the evolved $v_{\mathrm{l}}$, $v_{\mathrm{b}}$, and $z_{\mathrm{e}}$ in the same way as the calculation of the observed counterparts (see Section 2 and references therein). We overlaid the results of the evolved $v_{\mathrm{l}}$, $v_{\mathrm{b}}$, and $z_{\mathrm{e}}$ to Figure 2 in the same way as the initial counterparts. As is shown in Figure 2, the initial $v_{\mathrm{l}}$ distribution (labeled as “Tri-Normal”) is indeed in good consistency with the evolved $v_{\mathrm{l}}$ distribution, which validates the “stable-$v_{\mathrm{l}}$” assumption (see Section 3) within the initial phase space of this study. Despite the validation, we reiterate the necessity of assessing the consistency between the initial and the evolved $v_{\mathrm{l}}$ on a case-by-case basis, especially in the regime where $v_{\mathrm{sys}}$ and/or $\zeta_{0}$ are relatively large. In contrast, the distributions of evolved $v_{\mathrm{b}}$ and $z_{\mathrm{e}}$ are considerably different from the initial counterparts, which meets our expectation that the initial MSP $z_{\mathrm{i}}$ would spread out over time, while the initial $v_{\mathrm{b}}$ becoming weakened due to the pull of the Galactic potential.
Moreover, from the Galaxy-wide $z_{\mathrm{e}}$ sample (of $\sim 125$ million), we found that the best-fit scale height $\zeta_{1}$ for isolated MSP systems in the whole Galaxy is 0.66 kpc, which is 27% larger than the observed scale height $\zeta^{\mathrm{local}}_{\mathrm{obs}}=0.52$ kpc estimated from relatively nearby MSPs. Interestingly, unlike the observed $z$ sample that is best described by a Laplace distribution (see Appendix C.2.1), the Galaxy-wide $z_{\mathrm{e}}$ sample is best characterized by the Cauchy distribution with 0.47 kpc scale (see Figure 12).
5 The indicative millisecond pulsar retention fractions in star clusters
[…]
6 Summary and Prospects
This work derives the systemic recoil velocity ($v_{\mathrm{sys}}$) distribution of isolated millisecond pulsar (MSP) systems from the observed $v_{\mathrm{l}}$ (the Galactic-longitude component of the transverse peculiar velocity), assuming that 1) $v_{\mathrm{l}}$ is statistically stable over time (the “stable-$v_{\mathrm{l}}$” assumption), and 2) the $\bm{v}_{\mathrm{sys}}$ directions are uniformly distributed (the “isotropic-$\bm{v}_{\mathrm{sys}}$” assumption). We found that the $v_{\mathrm{sys}}$ distribution can be approximated by a linear combination of three Maxwellian distributions. Using dynamical population synthesis (DPS), we validated the “stable-$v_{\mathrm{l}}$” and evaluated the initial and current Galaxy-wide scale heights of isolated MSP systems by matching the observed Galactic heights and $v_{\mathrm{l}}$ (the Galactic-latitude component of the transverse peculiar velocity) to the DPS counterparts. The MSP $v_{\mathrm{sys}}$ distribution reported in this work predicts relatively large MSP retention fractions that may account for the large number of MSPs retained in globular clusters.
In light of the interesting discovery that the MSP $v_{\mathrm{sys}}$ distribution can be well approximated by a linear combination of three Maxwellian components, we highly encourage the production of template $v_{\mathrm{sys}}$ distribution for each formation channel using population synthesis that carefully simulates the MSP formation. These template $v_{\mathrm{sys}}$ distributions can be used to compare with the observation-based $v_{\mathrm{sys}}$ distribution, thereby deepening the understanding of MSP formation. Finally, the methods described in this paper can be used to determine the $v_{\mathrm{sys}}$ distribution and the scale height for other kinds of kinematically old compact stars (such as black hole systems) using precise astrometry (especially when radial velocity information is missing).
Acknowledgements.
HD appreciates the EACOA Fellowship awarded by the East Asia Core Observatories Association.
Appendix A Selection of the millisecond pulsar sample
We selected the MSP sample from the PSRCAT Catalogue1 (Manchester et al. 2005) with the help of the psrqpy Python package (Pitkin 2018). The MSPs were searched using the following criteria:
  1. (i)
    $P_{0}<\leavevmode\nobreak$60 ms, where $P_{0}$ represents the spin periods of MSPs, and
  2. (ii)
    $\dot{P}_{0}<2\times 10^{-17}\,{\mathrm{s\leavevmode\nobreak\ s^{-1}}}$, where $\dot{P}_{0}$ denotes the time derivative of $P_{0}$, and
  3. (iii)
    MSPs should not belong to globular clusters (as explained in Section 2), and
  4. (iv)
    both proper motion magnitude and parallax are of $>3\,\sigma$ significance, with the uncertainty in the proper motion magnitude calculated through the propagation of uncertainties in the right ascension (RA) and the declination components.
There is no clear $P_{0}$ boundary between MSPs and other pulsars. Various $P_{0}$ upper limits up to $\sim 100$ ms have been adopted in different studies of MSPs (e.g. Manchester 2017; Spiewak et al. 2022; Ding et al. 2023). In this work, we adopted the indicative $P_{0}$ upper limit of 60 ms. Additionally, we followed Spiewak et al. (2022), applying the filter $\dot{P}_{0}<2\times 10^{-17}\,\mathrm{s\leavevmode\nobreak\ s^{-1}}$ to exclude recently born ($\lesssim 10$ kyr old) pulsars (such as the Crab pulsar with 34 ms spin period), except for PSR J1417--4402 that does not have a reported $\dot{P}_{0}$. Despite the $\dot{P}_{0}$ being unknown to us, PSR J1417--4402 is suggested to have entered the late recycling phase (Camilo et al. 2016), therefore likely being kinematically similar to many other MSPs.
As mentioned in Section 2, we look for isolated MSP systems with $>3\,\sigma$ parallax determinations. However, the Gaia-based astrometric results registered in the PSRCAT Catalogue come from Jennings et al. (2018) based on the Gaia Data Release (DR) 2 (instead of the latest Gaia DR3). Therefore, in practice, we conducted the sample search in three steps. In the first step, we searched for MSPs with a discounted parallax significance threshold of $3\times 0.8=2.4$, where 0.8 is the expected ratio of parallax precision between the Gaia DR3 and the Gaia DR2. In the second step, we updated the astrometric results of Jennings et al. (2018) to Gaia DR3 utilizing the astroquery package (Ginsburg et al. 2019). In the final step, we removed MSPs with insignificant ($<3\,\sigma$) determinations of parallaxes or proper motions. Compared to parallaxes, proper motions are usually determined at higher significance. The proper motion filter (as part of the criterion iv) only removes one MSP (i.e., PSR J1811--2405) that is too close to the ecliptic plane to acquire precise proper motion in the declination direction with pulsar timing (Ng et al. 2020). Through the three-step procedure, we reached a sample of 65 isolated MSP systems, which are compiled in Table 2 along with their most precise proper motion and parallax determinations.
After calculating the transverse peculiar velocities $v_{\perp}$ of the 65 MSPs (see Section 2), we found no correlation between $P_{0}$ and $v_{\perp}$ (with Pearson coefficient of $-0.06$), hence indicating negligible selection effect of $P_{0}$ on $v_{\perp}$.
Table 2: Distances $d$, Galactic heights $z$, transverse peculiar velocities $\textit{{v}}_{\perp}\equiv\left(v_{\mathrm{l}},v_{\mathrm{b}}\right)$, and their magnitudes $v_{\perp}$ of the 65 selected millisecond pulsars calculated from their published proper motions ($\mu_{\alpha}$ and $\mu_{\delta}$) and parallaxes $\varpi$, where 1) negative $z$ values correspond to negative Galactic latitudes $b$, and 2) $v_{\mathrm{b}}$ and $v_{\mathrm{l}}$ represent the transverse peculiar velocity components along the directions of $b$ and Galactic longitudes $l$, respectively.
PSR $\mu_{\alpha}$ $\mu_{\delta}$ $\varpi$ Ref. $a$$a$$a$References for proper motions and parallaxes: (1) Bangale et al. (2024), (2) Alam et al. (2020), (3) Ding et al. (2023), (4) Shamohammadi et al. (2024), (5) Curyło et al. (2023), (6) Reardon et al. (2024), (7) van der Wateren et al. (2022), (8) Antoniadis et al. (2023b), (9) Guillemot et al. (2016), (10) Fiore et al. (2023), (11) Kramer et al. (2021), (12) Deller et al. (2019), (13) Deller et al. (2012), (14) Jennings et al. (2018), (15) Konacki & Wolszczan (2003), (16) Yan et al. (2013), (17) Swiggum et al. (2017), (18) Reardon et al. (2021), (19) Chatterjee et al. (2009), (20) Arzoumanian et al. (2018), (21) Deller et al. (2018), (22) Geyer et al. (2023), (23) Ding et al. (2024a), (24) Stovall et al. (2019), (25) Alam et al. (2021), and (26) Crowter et al. (2020).
$d$ $z$ $v_{l}$ $v_{b}$ $v_{\perp}$
($\mathrm{mas\leavevmode\nobreak\ yr^{-1}}$) ($\mathrm{mas\leavevmode\nobreak\ yr^{-1}}$) (mas) (kpc) (kpc) ($\mathrm{km\leavevmode\nobreak\ s^{-1}}$) ($\mathrm{km\leavevmode\nobreak\ s^{-1}}$) ($\mathrm{km\leavevmode\nobreak\ s^{-1}}$)
J0023$+$0923 $-11.00(7)$ $-8.8(1)$ $0.6(1)$ (2), (1) $1.9^{+0.5}_{-0.3}$ $-1.5^{+0.3}_{-0.4}$ $-89^{+16}_{-24}$ $-41^{+9}_{-13}$ $98^{+27}_{-19}$
J0030$+$0451 $-6.15(5)$ $0.4(1)$ $3.04(5)$ (3) $0.329(5)$ $-0.278(5)$ $-9.5(1)$ $12.2(2)$ $15.4(2)$
J0125--2327 $37.14(8)$ $10.5(2)$ $0.8(1)$ (5), (4) $1.2^{+0.2}_{-0.1}$ $-1.2^{+0.1}_{-0.2}$ $35^{+7}_{-5}$ $216^{+34}_{-25}$ $219^{+34}_{-26}$
J0437--4715 $121.4420(6)$ $-71.4717(7)$ $6.43(5)$ (6) $0.156(1)$ $-0.1040(8)$ $46.3(4)$ $85.3(7)$ $97.0(8)$
J0610--2100 $9.11(3)$ $16.45(3)$ $0.7(1)$ (7), (3) $1.5^{+0.3}_{-0.2}$ $-0.46^{+0.07}_{-0.09}$ $-72^{+10}_{-13}$ $95^{+19}_{-14}$ $119^{+24}_{-17}$
J0613--0200 $1.837(2)$ $-10.359(6)$ $1.00(5)$ (8) $1.01(5)$ $-0.163^{+0.008}_{-0.009}$ $44(3)$ $-17.7^{+0.8}_{-0.9}$ $47(3)$
J0614--3329 $0.58(9)$ $-1.9(1)$ $1.5(4)$ (4), (9) $1.0^{+0.7}_{-0.3}$ $-0.4^{+0.1}_{-0.3}$ $22^{+17}_{-8}$ $-8^{+2}_{-3}$ $23^{+18}_{-8}$
J0621$+$1002 $3.27(9)$ $-1.1(3)$ $0.7(1)$ (3) $1.6^{+0.5}_{-0.3}$ $-0.06^{+0.01}_{-0.02}$ $10^{+6}_{-4}$ $17^{+5}_{-4}$ $20^{+8}_{-5}$
J0636$+$5128 $1.1(4)$ $-4.4(7)$ $1.4(3)$ (10), (2) $0.9^{+0.3}_{-0.2}$ $0.28^{+0.11}_{-0.06}$ $8^{+8}_{-5}$ $-5^{+2}_{-3}$ $10^{+8}_{-4}$
J0636--3044 $26.01(9)$ $-19.3(1)$ $4.3(7)$ (4) $0.26^{+0.07}_{-0.04}$ $-0.07^{+0.01}_{-0.02}$ $34^{+10}_{-6}$ $17^{+5}_{-3}$ $39^{+11}_{-7}$
J0645$+$5158 $-7.44(4)$ $1.48(2)$ $0.8(2)$ (2) $1.4^{+0.4}_{-0.3}$ $0.48^{+0.15}_{-0.10}$ $-35^{+5}_{-8}$ $-47^{+9}_{-14}$ $59^{+16}_{-10}$
J0737--3039A $-2.57(3)$ $2.08(4)$ $1.3(1)$ (11) $0.80^{+0.09}_{-0.08}$ $-0.063^{+0.006}_{-0.007}$ $1.4^{+0.6}_{-0.5}$ $-6.2^{+0.5}_{-0.7}$ $6.4^{+0.8}_{-0.6}$
J0751$+$1807 $-2.70(1)$ $-13.27(7)$ $0.85(4)$ (8) $1.18^{+0.06}_{-0.05}$ $0.43(2)$ $55(3)$ $-37(2)$ $67^{+4}_{-3}$
J1012$+$5307 $2.624(3)$ $-25.487(4)$ $1.14(4)$ (3), (8) $0.88(3)$ $0.68(2)$ $82(3)$ $49(2)$ $95(4)$
J1022$+$1001 $-14.92(5)$ $5.61(4)$ $1.39(4)$ (12) $0.72(2)$ $0.56(2)$ $-42(1)$ $-21.8^{+0.8}_{-0.9}$ $47(1)$
J1023$+$0038 $4.76(3)$ $-17.34(4)$ $0.73(3)$ (13) $1.37^{+0.06}_{-0.05}$ $0.98(4)$ $122(5)$ $-33(2)$ $126^{+6}_{-5}$
J1024--0719 $-35.277(5)$ $-48.23(1)$ $0.93(5)$ (3), (8) $1.08(6)$ $0.70(4)$ $61^{+4}_{-3}$ $-291^{+15}_{-18}$ $297^{+17}_{-15}$
J1125--6014 $11.10(4)$ $-12.99(4)$ $1.2(3)$ (4) $2^{+5}_{-1}$ $0.03^{+0.08}_{-0.02}$ $214^{+498}_{-112}$ $-89^{+47}_{-209}$ $231^{+536}_{-121}$
J1227--4853 $-18.8(1)$ $7.30(9)$ $0.5(1)$ (14) $2.9^{+1.5}_{-0.8}$ $0.7^{+0.4}_{-0.2}$ $-183^{+50}_{-87}$ $71^{+37}_{-20}$ $196^{+97}_{-53}$
J1300$+$1240 $45.50(5)$ $-84.70(7)$ $1.41(8)$ (16), (15) $0.72(4)$ $0.69(4)$ $125^{+7}_{-6}$ $-300^{+17}_{-19}$ $325^{+20}_{-18}$
J1400--1431 $17(2)$ $-55(6)$ $4(1)$ (17) $0.7^{+1.2}_{-0.4}$ $0.5^{+0.9}_{-0.3}$ $-2^{+9}_{-23}$ $-199^{+104}_{-344}$ $204^{+351}_{-110}$
J1417--4402 $-4.76(4)$ $-5.10(5)$ $0.20(5)$ (14) $5.3^{+1.6}_{-1.0}$ $1.5^{+0.4}_{-0.3}$ $-14^{+18}_{-9}$ $-100^{+18}_{-23}$ $101^{+22}_{-15}$
J1431--4715 $-11.8(1)$ $-14.5(2)$ $0.5(1)$ (14) $2.6^{+1.4}_{-0.7}$ $0.5^{+0.3}_{-0.1}$ $-148^{+38}_{-67}$ $-114^{+30}_{-61}$ $189^{+95}_{-49}$
J1446--4701 $-4.23(10)$ $-2.4(2)$ $0.7(2)$ (4) $3^{+3}_{-1}$ $0.5^{+0.5}_{-0.2}$ $-7^{+23}_{-3}$ $-11^{+5}_{-11}$ $15^{+13}_{-6}$
J1455--3330 $7.85(1)$ $-1.98(4)$ $1.3(1)$ (8) $0.79^{+0.07}_{-0.06}$ $0.30^{+0.03}_{-0.02}$ $36(2)$ $-23(2)$ $43(3)$
J1518$+$4904 $-0.68(3)$ $-8.53(4)$ $1.24(4)$ (3) $0.81^{+0.03}_{-0.02}$ $0.66(2)$ $-13.4(5)$ $8.6(6)$ $16.0^{+0.8}_{-0.7}$
J1537$+$1155 $1.484(7)$ $-25.29(1)$ $1.07(7)$ (3) $0.95^{+0.07}_{-0.06}$ $0.71^{+0.05}_{-0.04}$ $-85^{+6}_{-7}$ $-56^{+3}_{-4}$ $102^{+8}_{-7}$
J1545--4550 $-0.50(6)$ $2.5(1)$ $0.4(1)$ (18) $5^{+4}_{-2}$ $0.6^{+0.4}_{-0.2}$ $112^{+199}_{-60}$ $40^{+35}_{-16}$ $119^{+204}_{-61}$
J1600--3053 $-0.943(3)$ $-6.92(1)$ $0.72(2)$ (8) $1.39(4)$ $0.39(1)$ $-22.5^{+0.8}_{-0.9}$ $-31.8^{+0.9}_{-1.0}$ $39(1)$
J1614--2230 $4.4(5)$ $-29(3)$ $1.5(1)$ (2) $0.66^{+0.05}_{-0.04}$ $0.23^{+0.02}_{-0.01}$ $-47(8)$ $-72(8)$ $86^{+11}_{-10}$
J1629--6902 $-6.50(3)$ $-8.44(4)$ $0.9(2)$ (4) $1.5^{+0.8}_{-0.4}$ $-0.36^{+0.09}_{-0.20}$ $-46^{+13}_{-25}$ $-1.6(3)$ $46^{+26}_{-13}$
J1640$+$2224 $2.102(4)$ $-11.333(7)$ $0.73(6)$ (3), (8) $1.4(1)$ $0.86^{+0.08}_{-0.07}$ $-43(4)$ $-30(2)$ $53^{+5}_{-4}$
J1643--1224 $6.1(4)$ $3(2)$ $(1)1$ (3) $0.94^{+0.10}_{-0.08}$ $0.34^{+0.04}_{-0.03}$ $38^{+8}_{-7}$ $-13(5)$ $40^{+6}_{-5}$
J1658--5324 $0.0(1)$ $2.9(2)$ $1.3(3)$ (4) $1.4^{+3.2}_{-0.5}$ $-0.16^{+0.06}_{-0.37}$ $33^{+98}_{-9}$ $13^{+33}_{-5}$ $35^{+105}_{-11}$
J1713$+$0747 $4.9215(8)$ $-3.920(2)$ $0.95(6)$ (8), (19) $1.07^{+0.07}_{-0.06}$ $0.45(3)$ $9.8(1)$ $-28(2)$ $29^{+2}_{-1}$
J1723--2837 $-11.73(4)$ $-24.05(3)$ $1.07(4)$ (14) $0.94(4)$ $0.070(3)$ $-109(5)$ $-17.0^{+0.7}_{-0.8}$ $110(5)$
J1730--2304 $20.1(1)$ $-4.7(6)$ $2.0(1)$ (3) $0.51^{+0.03}_{-0.02}$ $0.053(3)$ $27(1)$ $-47^{+2}_{-3}$ $54^{+3}_{-2}$
J1738$+$0333 $7.09(1)$ $5.07(3)$ $0.59(5)$ (3), (8) $1.7^{+0.2}_{-0.1}$ $0.53^{+0.05}_{-0.04}$ $87^{+8}_{-7}$ $-27(2)$ $91^{+8}_{-7}$
J1741$+$1351 $-8.98(2)$ $-7.41(2)$ $0.6(1)$ (20) $1.8^{+0.4}_{-0.3}$ $0.7^{+0.2}_{-0.1}$ $-59^{+10}_{-13}$ $53^{+13}_{-9}$ $79^{+19}_{-13}$
J1744--1134 $18.806(2)$ $-9.39(1)$ $2.58(3)$ (8) $0.388(5)$ $0.0618(7)$ $13.43(4)$ $-38.5(4)$ $40.8(4)$
J1757--5322 $-2.48(8)$ $-10.0(2)$ $1.2(3)$ (4) $1.4^{+1.9}_{-0.5}$ $-0.3^{+0.1}_{-0.4}$ $-52^{+22}_{-67}$ $-13^{+4}_{-13}$ $53^{+71}_{-22}$
J1804--2717 $2.46(2)$ $-16.9(4)$ $1.1(3)$ (8) $6^{+7}_{-4}$ $-0.3^{+0.2}_{-0.3}$ $-252^{+148}_{-94}$ $-284^{+195}_{-347}$ $412^{+310}_{-278}$
J1832--0836 $-8.06(5)$ $-21.0(2)$ $0.5(1)$ (2), (18) $7^{+5}_{-4}$ $0.03(2)$ $-546^{+229}_{-320}$ $-84^{+43}_{-62}$ $551^{+304}_{-232}$
J1853$+$1303 $-1.62(2)$ $-2.96(4)$ $0.53(7)$ (3) $2.0^{+0.4}_{-0.3}$ $0.19^{+0.03}_{-0.02}$ $13^{+3}_{-2}$ $3.5^{+0.8}_{-0.6}$ $14^{+3}_{-2}$
J1857$+$0943 $-2.670(3)$ $-5.428(6)$ $0.89(6)$ (8) $1.15^{+0.09}_{-0.07}$ $0.061^{+0.005}_{-0.004}$ $-7.7^{+0.8}_{-0.9}$ $0.11(4)$ $7.7(9)$
J1909--3744 $-9.523(1)$ $-35.780(5)$ $0.94(2)$ (8) $1.07(2)$ $-0.357^{+0.007}_{-0.008}$ $-176(4)$ $-15.5(3)$ $177(4)$
J1910$+$1256 $0.24(2)$ $-7.10(3)$ $0.36(6)$ (3), (8) $3.2^{+0.9}_{-0.6}$ $0.10^{+0.03}_{-0.02}$ $-12.7^{+4.1}_{-0.9}$ $-52^{+9}_{-14}$ $53^{+13}_{-8}$
J1911$+$1347 $-2.900(5)$ $-3.684(9)$ $0.40(9)$ (8) $3.5^{+2.1}_{-1.0}$ $0.11^{+0.07}_{-0.03}$ $16^{+32}_{-9}$ $16^{+9}_{-4}$ $23^{+31}_{-10}$
J1915$+$1606 $-0.7(1)$ $-0.0(1)$ $0.24(7)$ (21) $6^{+3}_{-2}$ $0.23^{+0.11}_{-0.07}$ $185^{+87}_{-67}$ $20^{+9}_{-6}$ $185^{+86}_{-65}$
J1918--0642 $-7.124(5)$ $-5.96(2)$ $0.71(7)$ (3), (8) $1.5^{+0.2}_{-0.1}$ $-0.23^{+0.02}_{-0.03}$ $-37^{+4}_{-5}$ $23(2)$ $44^{+6}_{-4}$
J1933--6211 $-5.62(1)$ $11.09(3)$ $0.7(2)$ (22), (4) $1.9^{+1.1}_{-0.5}$ $-0.9^{+0.3}_{-0.5}$ $118^{+68}_{-31}$ $68^{+41}_{-20}$ $136^{+80}_{-37}$
\@makecaption
continued
PSR $\mu_{\alpha}$ $\mu_{\delta}$ $\varpi$ Ref. $a$$a$$a$References for proper motions and parallaxes: (1) Bangale et al. (2024), (2) Alam et al. (2020), (3) Ding et al. (2023), (4) Shamohammadi et al. (2024), (5) Curyło et al. (2023), (6) Reardon et al. (2024), (7) van der Wateren et al. (2022), (8) Antoniadis et al. (2023b), (9) Guillemot et al. (2016), (10) Fiore et al. (2023), (11) Kramer et al. (2021), (12) Deller et al. (2019), (13) Deller et al. (2012), (14) Jennings et al. (2018), (15) Konacki & Wolszczan (2003), (16) Yan et al. (2013), (17) Swiggum et al. (2017), (18) Reardon et al. (2021), (19) Chatterjee et al. (2009), (20) Arzoumanian et al. (2018), (21) Deller et al. (2018), (22) Geyer et al. (2023), (23) Ding et al. (2024a), (24) Stovall et al. (2019), (25) Alam et al. (2021), and (26) Crowter et al. (2020).
$d$ $z$ $v_{l}$ $v_{b}$ $v_{\perp}$
($\mathrm{mas\leavevmode\nobreak\ yr^{-1}}$) ($\mathrm{mas\leavevmode\nobreak\ yr^{-1}}$) (mas) (kpc) (kpc) ($\mathrm{km\leavevmode\nobreak\ s^{-1}}$) ($\mathrm{km\leavevmode\nobreak\ s^{-1}}$) ($\mathrm{km\leavevmode\nobreak\ s^{-1}}$)
J1939$+$2134 $0.074(2)$ $-0.410(3)$ $0.35(3)$ (3) $3.0^{+0.3}_{-0.2}$ $-0.015(1)$ $80^{+9}_{-7}$ $-3.8^{+0.3}_{-0.4}$ $80^{+9}_{-8}$
J1946--5403 $-1.08(2)$ $-4.75(5)$ $0.8(1)$ (4) $1.3(2)$ $-0.65^{+0.08}_{-0.11}$ $-18^{+4}_{-5}$ $7^{+2}_{-1}$ $19^{+5}_{-4}$
J2043$+$1711 $-5.70(1)$ $-10.84(2)$ $0.72(7)$ (2) $1.4^{+0.2}_{-0.1}$ $-0.38(4)$ $-43^{+4}_{-5}$ $-13^{+1}_{-2}$ $45^{+5}_{-4}$
J2124--3358 $-14.09(1)$ $-50.32(3)$ $2.1(1)$ (8) $0.48(2)$ $-0.34(2)$ $-106^{+5}_{-6}$ $20.6^{+1.0}_{-0.9}$ $108^{+6}_{-5}$
J2129--0429 $12.10(7)$ $10.19(6)$ $0.48(7)$ (1), (14) $2.1^{+0.4}_{-0.3}$ $-1.3(2)$ $194^{+32}_{-25}$ $-63^{+9}_{-11}$ $204^{+37}_{-27}$
J2145--0750 $-9.7(3)$ $-8.2(9)$ $1.60(6)$ (12) $0.63(2)$ $-0.42(2)$ $-21(3)$ $14(2)$ $25(2)$
J2222--0137 $44.707(5)$ $-5.40(1)$ $3.72(1)$ (23) $0.269(1)$ $-0.1935(7)$ $39.8(1)$ $-43.3(2)$ $58.8(2)$
J2234$+$0611 $25.30(2)$ $9.71(5)$ $1.03(4)$ (24) $0.97(4)$ $-0.66(3)$ $139^{+6}_{-5}$ $-44(2)$ $145^{+6}_{-5}$
J2234$+$0944 $7.0(2)$ $-33.2(5)$ $1.4(3)$ (25) $0.9^{+0.3}_{-0.2}$ $-0.6^{+0.1}_{-0.2}$ $-46^{+11}_{-18}$ $-116^{+26}_{-45}$ $124^{+49}_{-28}$
J2241--5236 $18.87(1)$ $-5.28(1)$ $0.96(4)$ (18) $1.05^{+0.05}_{-0.04}$ $-0.86^{+0.03}_{-0.04}$ $-62(3)$ $-61^{+2}_{-3}$ $87(4)$
J2256--1024 $3(1)$ $-8(2)$ $0.5(2)$ (26) $2.1^{+0.7}_{-0.5}$ $-1.8^{+0.4}_{-0.6}$ $-18^{+17}_{-21}$ $-79^{+22}_{-35}$ $83^{+35}_{-25}$
J2317$+$1439 $-1.46(10)$ $3.7(2)$ $0.60(8)$ (2) $1.7^{+0.3}_{-0.2}$ $-1.2^{+0.1}_{-0.2}$ $40^{+6}_{-5}$ $42^{+6}_{-5}$ $58^{+9}_{-7}$
J2322$+$2057 $-18.30(5)$ $-14.9(1)$ $1.2(3)$ (8), (4) $1.0^{+0.5}_{-0.2}$ $-0.6^{+0.2}_{-0.3}$ $-91^{+22}_{-43}$ $-19^{+6}_{-12}$ $93^{+43}_{-23}$
J2322--2650 $-2.37(9)$ $-8.2(2)$ $1.3(2)$ (4) $0.8^{+0.2}_{-0.1}$ $-0.8^{+0.1}_{-0.2}$ $-21^{+4}_{-6}$ $7.4^{+0.6}_{-0.5}$ $23^{+6}_{-4}$
333 When an MSP has multiple proper motions and parallaxes reported by different studies, the values with smaller uncertainties are adopted. Despite our meticulous efforts, some of the compiled proper motions and parallaxes may not represent the most precise measurements available.
Appendix B Determining the probability density function for $v_{\mathrm{l}}$
B.1 Candidate probability density functions of $v_{\mathrm{l}}$
Following the first and the second steps described in Section 3, we identified two prime groups of candidate probability density functions (PDFs) of $v_{\mathrm{l}}$. The first group is a linear combination of Cauchy distributions where the probability density $p$ takes the form of
$\cases{p}=\displaystyle\sum_{i}^{N}\frac{\alpha_{i}}{\pi\gamma_{i}\left(1+\frac{{v_{\mathrm{l}}}^{2}}{{\gamma_{i}}^{2}}\right)}\,\,\,\,\,\,\,\,\,\,(\alpha_{i}>0\text{ and }\gamma_{i}>0)\\ \displaystyle\sum_{i}^{N}\alpha_{i}=1\,\,\,,{}$ (2)
while the second group being a linear combination of normal distributions, following the relation
[…]
In this paper, we refer to Equations 2 and 3 as Cauchy series and Gaussian series, respectively. In either equation, $\gamma_{i}$ is the scale of the $i$-th Cauchy or normal distribution. When $N=1$, Equations 2 and 3 are simplified to a Cauchy distribution and a normal distribution, respectively. For both groups, we refer to a PDF with $N=2,3,4$ as a two-component distribution, a three-component distribution, and a four-component distribution, respectively. The best-fit parameters for Equations 2 and 3 are provided in Table 4.
Table 4: Best-fit parameters of candidate $v_{\mathrm{l}}$ probability density functions defined by Equations 2 and 3, along with the Bayes factors $K$ and the goodness differences $\Delta\Lambda$ calculated with the best-fit Cauchy distribution ($N=1$ and $\gamma_{1}=51.8$$\mathrm{km\leavevmode\nobreak\ s^{-1}}$) as the reference of comparison.
$N$ $\gamma_{1}$ $\alpha_{1}$ $\gamma_{2}$ $\alpha_{2}$ $\gamma_{3}$ $\alpha_{3}$ $\gamma_{4}$ $\alpha_{4}$ $\ln{K}$ $\Delta\Lambda$
($\mathrm{km\leavevmode\nobreak\ s^{-1}}$) ($\mathrm{km\leavevmode\nobreak\ s^{-1}}$) ($\mathrm{km\leavevmode\nobreak\ s^{-1}}$) ($\mathrm{km\leavevmode\nobreak\ s^{-1}}$)
Cauchy series (Eq. 2)
1 51.8 100% 0 0
2 51.8 99.2% 52.4 0.8% $6^{+3}_{-2}\!\times\!10^{-4}$ $26^{+3}_{-2}\!\times\!10^{-4}\!-\!$ 2
Gaussian series (Eq. 3)
1 68.6 100% $-14^{+8}_{-13}$ $-14^{+8}_{-13}$
2 46.8 54.2% 127.4 45.8% $-2^{+2}_{-4}$ $-4^{+2}_{-4}$
3 45.5 49.8% 110.3 44.7% 448.8 5.5% 0.6(2) $-3.4(2)$
4 45.5 49.7% 110.2 44.8% 391.5 1.8% 486.8 3.7% 0.6(2) $-5.4(2)$
B.2 Selecting from the candidate probability density distributions
From each of the two prime PDF groups introduced in Appendix B.1, we selected one $v_{\mathrm{l}}$ PDF based on a criterion that is equivalent to the Akaike information criterion (Akaike 2011). Specifically, the goodness $\Lambda$ of a PDF is quantified by
$\Lambda=\ln{L}-k\,\,,$ (4)
where $L$ refers to the likelihood that the measurements match the PDF, and $k$ represents the number of free parameters in the PDF. A PDF with higher $\Lambda$ is considered better. For Equations 2 and 3,
$k=2N-1\,\,;$ (5)
the likelihood $L$ can be calculated by
$L=\displaystyle\prod_{j=1}^{65}p\left(v^{j}_{\mathrm{l}}\right)\,\,,$ (6)
where $p\left(v^{j}_{\mathrm{l}}\right)$ is the value of Equation 2 or 3 at the observed $v^{j}_{\mathrm{l}}\,\,(j=1,2,...,65)$.
Accordingly, the comparison of two PDFs is based on the $\Lambda$ difference
$\Delta\Lambda=\Lambda_{1}-\Lambda_{2}=\ln{\left(\frac{L_{1}}{L_{2}}\right)}-(k_{1}-k_{2})=\ln{K}-2\left(N_{1}-N_{2}\right)\,\,,$ (7)
where $K\equiv L_{1}/L_{2}$ is the Bayes factor between the two PDFs. To evaluate the uncertainty in $K$ due to the uncertainties in the 65 $v_{\mathrm{l}}$ measurements, we assumed the observed $v_{\mathrm{l}}$ follow split normal distributions, and simulated 10000 sets of 65 $v_{\mathrm{l}}$ values. From each set of 65 $v_{\mathrm{l}}$ values we derived a $K$ based on Equation 6. With the altogether 10000 derived $K$ values, we estimated $K$ and its uncertainty for two PDFs of comparison. The calculated $\ln{K}$ (as well as their uncertainties) of the candidate PDFs with respect to the best-fit Cauchy distribution are provided in Table 4.
With larger $N$ (and accordingly more free parameters), the PDF is expected to better approach the data, leading to larger $K$. Meanwhile, the incremental increase of $\ln{K}$ would quickly drop at larger $N$. For each of the two prime groups of candidate PDFs, we started fitting the PDF parameter(s) and estimating the corresponding $K$ from $N=1$, then repeated the parameter fitting and $K$ estimation at incrementally larger $N$ until the incremental increase of $\ln{K}$ becomes negligible compared to 2 – the incremental increase of the number of free parameters. We found that $\Lambda$ (defined by Equations 4 through 6) peaks at $N=1$ for the Cauchy series, and at $N=3$ for the Gaussian series (see Table 4). In other words, the best-fit Cauchy distribution and the best-fit three-component normal distribution are selected for the two prime groups of candidate PDFs.
Between the two finalist PDFs, the Cauchy distribution is preferred over the three-component normal distribution according to the $\Lambda$ standard (see Table 4). Therefore, we conclude that the sample of 65 observed $v_{\mathrm{l}}$ is well described by the simple PDF
$\frac{p}{1\,\mathrm{s\leavevmode\nobreak\ {km}^{-1}}}=\frac{1}{51.8\pi\left[1+\left(\frac{{v_{\mathrm{l}}}}{51.8\,\mathrm{km\leavevmode\nobreak\ s^{-1}}}\right)^{2}\right]}\,\,.$ (8)
On the other hand, the three-component normal distribution is highly convenient for the calculation of the $v_{\mathrm{sys}}$ distribution: when assuming $\bm{v}_{\mathrm{sys}}$ is statistically isotropic (in other words, equally possible in all directions) for kinematically old pulsar systems, the $\bm{v}_{\mathrm{sys}}$ magnitude $v_{\mathrm{sys}}$ would follow a three-component Maxwell distribution with parameters identical to those of the three-component normal distribution (see Equation 1). Furthermore, the analytical solution
$\mathcal{F}\left(v_{\mathrm{sys}}\right)=\sum_{i=1}^{3}\alpha_{i}\left[\mathrm{erf}\left(\frac{v_{\mathrm{sys}}}{\sqrt{2}\gamma_{i}}\right)-\sqrt{\frac{2}{\pi}}\cdot\frac{v_{\mathrm{sys}}}{\gamma_{i}}\exp{\left(-\frac{{v_{\mathrm{sys}}}^{2}}{2\gamma_{i}^{2}}\right)}\right]$ (9)
can be derived for the cumulative distribution function (CDF) of the three-component Maxwell distribution, where erf is the error function, and the parameters $\alpha_{i}$ and $\gamma_{i}$ can be found in Table 4. The three-component Maxwell distribution as well as its corresponding CDF is displayed in Figure 4. In comparison, provided the same “isotropic-$\bm{v}_{\mathrm{sys}}$” assumption, no closed-form solution of $v_{\mathrm{sys}}$ can be obtained when each direction of $\bm{v}_{\mathrm{sys}}$ follows a Cauchy distribution. Therefore, we consider the three-component normal distribution a good and useful approximation to Equation 8 and the $v_{\mathrm{l}}$ PDF.
Figure 4: Upper: The three-component Maxwell distribution described by Equation 1, which shares the parameters of the best-fit three-component normal distribution (see Table 4) of the observed $v_{\mathrm{l}}$ sample. Lower: The cumulative distribution function (CDF) calculated from the three-component Maxwell distribution shown in the upper panel. The dash-dotted vertical line marks the mode of the three-component Maxwell distribution, while the dashed lines and the dotted line correspond to the 16th, the 84th, and the 50th percentile of the CDF, respectively.
Appendix C Supplementary materials on dynamical population synthesis
C.1 Millisecond pulsar age distribution
The age distribution of MSPs is poorly constrained. We approached the rough MSP age distribution by the following procedure. Firstly, we identified MSPs from the PSRCAT1 catalogue with the MSP criteria (i) and (ii) stated in Appendix A. After removing sources without constraint on characteristic age $\tau_{\mathrm{c}}$, we obtained 399 MSP $\tau_{\mathrm{c}}$ values. To derive the age distribution, we corrected the $\tau_{\mathrm{c}}$ sample by a sample of age correction factors (i.e., ratio between age and $\tau_{\mathrm{c}}$) estimated by Kiziltan & Thorsett (2010) for 42 MSPs having proper motion measurements (i.e., the reciprocals of “$\tau_{c}/\tilde{\tau}^{\prime}$” reported in Table 1 of Kiziltan & Thorsett 2010). Specifically, we multiplied each of the $\tau_{\mathrm{c}}$ sample by each of the correction factor sample, which led to a sample of $399\times 42=16,758$ MSP ages. After discarding ages larger than the age of the Galaxy (i.e., 13.6 Gyr), we obtained a sample of 13,005 MSP ages (see Figure 6 for the histogram of this sample). We considered this age sample as an indicative approximation to the MSP age distribution, and adopted the age sample in the DPS analysis (see Section 4.2).
Despite the yet poorly constrained MSP age distribution, we find that the DPS results are insensitive to the uncertainties in the age distribution (see Appendix C.1.1). While this insensitivity reinforces the robustness of the DPS results, it also prohibits the possibility of constraining age distribution of MSPs (and likely other kinematically old compact star systems) through astrometric studies.
Figure 6: Gray: The histogram of the millisecond pulsar age distribution adopted in DPS (see Appendix C.1), with logarithmically spaced bins along the x-axis; Pink: The uniform distribution $\mathcal{U}\left(0.1,13.6\right)$ (Gyr) with logarithmically spaced bins along the x-axis, which is only used to examine the DPS robustness with respect to uncertainties in the MSP age distribution (see Appendix C.1.1).
C.1.1 The robustness with respect to uncertainties in the age distribution
To investigate the impact of the uncertainties in the MSP age distribution on the DPS results, we reran DPS with the same setup except for the MSP age distribution. For the rerun, we adopted the uniform distribution $\mathcal{U}\left(0.1,13.6\right)$ (Gyr) as the MSP age distribution (i.e., the ages are evenly distributed between 0.1 Gyr and 13.6 Gyr), which is overlaid to Figure 6 in order to show its deviation from the age distribution of reference. Despite the difference between the uniform distribution and the reference distribution, we find that the two distinct age distributions render almost identical DPS results (see Figure 8). Therefore, we conclude that the DPS results are robust with respect to the uncertainties in the MSP age distribution.
Figure 8: Difference between the smoothed histogram profiles of evolved $v_{\mathrm{l}}$, $v_{\mathrm{b}}$, and $z_{\mathrm{e}}$ based on two distinct MSP age distributions — the uniform distribution $\mathcal{U}\left(0.1,13.6\right)$ (Gyr) in blue, and the MSP age distribution described in Appendix C.1.
C.2 Supplementary materials on the determination of the birth Galactic height distribution
C.2.1 The underlying probability density function of the observed Galactic heights
When searching for the candidates for the underlying PDF of the observed $z$ sample, we assumed that the PDF is both uni-modal and symmetric about zero. The same assumptions had been adopted in identifying the underlying PDF for the observed $v_{\mathrm{l}}$ (see Section 3). We found that the one-parameter Cauchy distribution and the one-parameter Laplace distribution are both plausible candidates for the underlying PDF of the observed $z$ (see Figure 2). However, further analysis following the method detailed in Appendix B.2 prefers the best-fit Laplace distribution over the Cauchy counterpart with $\ln{K}=8.0^{+0.7}_{-0.9}$, where $K$ is the Bayes factor. The best-fit Laplace distribution for the observed $z$ sample has the scale height $\zeta^{\mathrm{local}}_{\mathrm{obs}}$ of 0.52 kpc. We note that $\zeta^{\mathrm{local}}_{\mathrm{obs}}$ is derived from nearby isolated MSP system, which should be distinguished from the scale height $\zeta_{1}$ of Galaxy-wide isolated MSP systems.
C.2.2 The determination of the Galaxy-wide birth scale height $\zeta_{0}$
As mentioned in Section 4.1, various $\zeta_{0}$ would lead to different levels of the evolved $v_{\mathrm{b}}$ and $z_{\mathrm{e}}$. Due to the complex distributions of the observed $z$ and $v_{\mathrm{b}}$ (see Figure 2 or Figure 10), we find it difficult to determine the $\zeta_{0}$ using statistical methods (e.g., Kolmogorov-Smirnov tests or the Wasserstein metric) that quantify the similarity between the observed samples and the DPS counterparts. On the other hand, the evolved scale height $\zeta_{\mathrm{e}}^{\mathrm{local}}$ is expected to increase with $\zeta_{0}$ (see Figure 10), and match the observed counterpart $\zeta^{\mathrm{local}}_{\mathrm{obs}}=0.52$ kpc (see Appendix C.2.1). Therefore, we reran the DPS with incrementally larger $\zeta_{0}$ until the best-fit $\zeta_{\mathrm{e}}^{\mathrm{local}}$ reaches 0.52 kpc. In this way, we found $\zeta_{0}\approx 0.32$ kpc. The dependence of the evolved $v_{\mathrm{b}}$, $z_{\mathrm{e}}$ and $v_{\mathrm{l}}$ on $\zeta_{0}$ is illustrated in Figure 10. As is shown in Figure 10, both the evolved $v_{\mathrm{b}}$ and $z_{\mathrm{e}}$ match the observed counterparts at $\zeta_{0}=0.32$ kpc. Additionally, the average levels of both evolved $|\mathrm{b}|$ and $|z_{\mathrm{e}}|$ rise with $\zeta_{0}$ as expected. In comparison, the evolved $v_{\mathrm{l}}$ stays robust with respect to changes in the birth vertical distribution of isolated MSP systems.
Figure 10: The smoothed histogram profiles of $z$, $v_{\mathrm{l}}$, and $v_{\mathrm{b}}$ derived from dynamical population synthesis based on different Galaxy-wide birth scale height $\zeta_{0}$ of isolated MSP systems. Overlaid are the visualized histograms (and their smoothed profiles) of the observed $z$, $v_{\mathrm{l}}$, and $v_{\mathrm{b}}$ (see Section 2).
Finally, we note that the observed $z$ sample might be biased by Galactic-latitude-related selection effects. As a major example, pulsars located at low Galactic latitudes ($b$) are more susceptible to propagation effects caused by the ionized interstellar medium (IISM), which might reduce the detection rate of these low-$b$ pulsars. This relative drop in the number of low-$b$ pulsars would become increasingly significant as further-away pulsars are searched and observed with large telescopes, but may still have small impact on this work where the vast majority of MSPs are in the vicinity of the Solar system (see Table 2). On the other hand, most pulsar search programs have been focusing on low-$b$ sky regions, which serves as another selection effect that would increase the relative number of low-$b$ pulsars. For this work, we see no need to take into account the impact of complex $b$-related selection effects on the observed $z$, especially as the DPS results based on $\zeta_{0}=0.32$ kpc agree with both the observed $z$ and the observed $v_{\mathrm{b}}$ samples. However, if the observed $z$ and the observed $v_{\mathrm{b}}$ suggest different $\zeta_{0}$, this may indicate that $b$-related selection effects have to be accounted for.
Figure 12: The histogram of Galaxy-wide evolved Galactic heights $z_{\mathrm{e}}$ derived from dynamical population synthesis. Overlaid are the best-fit Laplace and Cauchy distributions of the $z_{\mathrm{e}}$ sample, with the scale parameters of 0.66 kpc and 0.47 kpc, respectively.